OpenStudy (anonymous):

Is log 5 (lower) 125^x simplified as 3x?

OpenStudy (anonymous):

yes because \[125=5^3\]so \[log_5(125^x)=log_5((5^3)^x)=log_5(5^{3x})=3x\]

OpenStudy (anonymous):

thanx i was just checking